# Uncovering Bias in Large Vision-Language Models at Scale with Counterfactuals

### v1

@@ Appendix A Limitations and Ethical Considerations
We do not foresee any negative societal impacts resulting from the work performed during this study. On the contrary, we believe the findings from this paper will raise awareness of potential risks and harms when LVLMs are deployed at scale. We hope that our work encourages future research aimed at reducing such risks, and believe it will consequently have positive societal impacts through the development of more fair and responsible AI models. Without awareness of the limitations, bias, stereotypes and toxicity of LVLMs, we risk not just correctness but also fairness to demographic groups, genders and occupations. With some of the mitigation strategies we present in the study, we intend to inspire further innovation in improving LVLMs.
Appendix B SocialCounterfactuals Dataset
@@
The SocialCounterfactuals dataset was introduced by Howard et al. [9] and generated automatically using Stable Diffusion with cross-attention control in order to produce visually consistent counterfactual sets. Each set portrays a person in an occupation-based scenario (accountant, pastry chef, plumber, surgeon, etc); however, each image in the set varies in terms of the social attributes of the depicted person. There are three intersectional subsets: Race-Gender, Physical-Gender, and Physical-Race. The size of each subset is given in Table 8. The attribute Gender can have one of two values, male or female. The attribute Race can have one of six values: White, Black, Indian, Asian, Middle Eastern, or Latino. The Physical Characteristics attribute can have one of five values: skinny, obese, young, old, or tattooed.
Subset Counterfactual Sets Images Per Set Total Images
@@
Table 8: Number of images in each of the three intersectional subsets of SocialCounterfactuals (adapted from [9]).
Appendix C Differences in Model Response Length across Social Groups
In our experiments, we generate responses of up to 512 tokens from each LVLMs. The vast majority of model responses are much shorter than this limit because an end-of-sequence or end-of-turn token terminates generation early. This allows us to study how intersectional social groups depicted in an image influence the amount of text that LVLMs generate in response to different prompts.
@@ Appendix C Differences in Model Response Length across Social Groups
Appendix D Additional Results and Details from LVLM Toxicity Analysis
D.1 Toxicity Evaluation with Real Image Datasets
Our use of the synthetic SocialCounterfactual dataset for bias evaluations may raise the question of the extent to which similar biases are observed when LVLMs are presented with real images. Unfortunately there are no natural counterfactual image datasets which cover intersectional social attributes at the scale of SocialCounterfactuals. The Protected-Attribute Tag Association (PATA) [22] dataset contains 4,934 images organized in 24 scenes with binary genders (female, male), five ethnic-racial labels (Asian, Black, Hispanic, Indian, White) and two age group labels (young and old). We attempt to align the attributes from PATA to those defined in SocialCounterfactuals.
We evaluate the same five open LVLMs on this dataset using identical prompts as our previous experiments, varying the random seed 15 times per image to produce a comparable number of model responses. While we cannot calculate MaxToxicity because PATA lacks counterfactual sets, we report the 90th percentile of Perspective API toxicity scores for each model. Table 9 provides these values by intersectional race-gender groups which have been aligned to the same labels as SocialCounterfactuals. We observe similar bias trends as was described previously; specifically, BakLLaVA consistently exhibits the highest toxicity across the five models, with all models producing the most toxic content for images depicting Black subjects. While these results are not directly comparable to those obtained using the SocialCounterfactual datasets due to the lack of counterfactual evaluation sets in the PATA dataset, they indicate that bias in the generation of toxic content is not solely a consequence of using synthetic images as input.
@@ D.1 Toxicity Evaluation with Real Image Datasets
D.2 Variation in Toxicity Scores by Occupation Depicted in the Image
@@
D.3 MaxToxicity Evaluation with Other Prompts
@@
D.4 Evaluation of Open-Source LVLMs with Other Perspective API Scores
@@
$\textrm{MaxInsult}_{c}=\max_{(a_{i},a_{j})\in A}\bigr{[}I(x)_{c,a_{i},a_{j}}\bigr{]}-\min_{(a_{i},a_{j})\in A}\bigr{[}I(x)_{c,a_{i},a_{j}}\bigr{]}\$ (2)
@@
$\textrm{MaxFlirtation}_{c}=\max_{(a_{i},a_{j})\in A}\bigr{[}F(x)_{c,a_{i},a_{j}}\bigr{]}-\min_{(a_{i},a_{j})\in A}\bigr{[}F(x)_{c,a_{i},a_{j}}\bigr{]}\$ (4)
@@
D.5 Analysis of BakLLaVA Flirtation Scores by Physical-Gender Attributes and Occupation
@@
To better understand the high values of flirtation scores for BakLLaVA responses to the characteristics prompt (Table 13), we analyze a subset of these generations for 8 occupations which had the highest standard deviation of flirtation scores across physical-gender groups. Figure 8 provides boxplots of BakLLaVA’s Flirtation scores for these response, broken down by intersectional physical-gender groups. While we observe higher flirtation scores for female subjects in general, skinny, young, and tattooed females have particularly high Flirtation scores relative to other groups across occupations such as Dentist, Bartender, Cashier, and Driver. The high degree of variability across different occupations suggests that bias in the generation of flirtatious content is highly influenced by the occupation depicted in the image.
D.6 Inference-Time Mitigation
@@
Table 14 provides details of the five mitigation instructions that we investigated. Mitigation instructions are added either before or after our original prompts, which are represented with the {Prompt} placeholder. In addition to evaluating mitigation effectiveness for physical-gender images and the Characteristics prompt (Table 7), we also evaluated the effectiveness of the M1 mitigation instruction using all three intersectional attribute types and other prompts used throughout our analyses. Tables 15 and 16 provide the reduction in MaxToxicity scores when the M1 mitigation instruction is used across these various evaluation settings. We find that this instruction is most effective for the Personality and Trustworthiness prompts, particularly for images depicting intersectional physical-gender social attributes. However, the high degree of variability across prompts (even for the same model and social attribute types) indicates that a single inference-time mitigation instruction is unlikely to reduce bias across a broad-range of generation scenarios. In many cases (e.g., BakLLaVA physical-gender and physical-race evaluations), large reductions in MaxToxicity for the Personality, Characteristics, and Trustworthiness prompts are contrasted with increases in MaxToxicity for the Describe, Backstory, and Pretend prompts. This highlights the need for additional research into robust methodologies for reducing bias in LVLMs.
D.7 LLM-LVLM Mapping
In Section 5.1, we analyze the relationship between MaxToxicity observed in LVLMs and that of the LLM from which they was trained. Table 17 provides a mapping between the LVLMs used in our study and the LLMs from which they were derived. We use these corresponding LVLM & LLM pairs for the purpose of our analyses in Section 5.1.
LVLM LLM
LLaVA-7B Vicuna-7b-v1.5
LLaVA-13B Vicuna-13b-v1.5
BakLLaVA Mistral-7B-Instruct-v0.1
InstructBLIP Vicuna-7b-v1.5
LLaVA-Gemma Gemma-2b-it
Table 17: Mapping of LVLMs to the LLM from which they were trained
Appendix E Additional Examples
Appendix F Generation Details
F.1 Text Prompts
Table 18 provides the open-ended text prompts that were investigated in this study, which were originally proposed by  Fraser and Kiritchenko [5]. We refer to each prompt throughout this paper by its short-hand label indicated in Table 18. Table 19 provides additional prompts that we investigated; we provide results for these additional prompts only in the subsequent sections of this Appendix.
Label Prompt
Describe Describe the image in as much detail as possible.
Table 18: Text prompts used for main analyses.
Label Prompt
Table 19: Additional prompts investigated during this study.
F.2 Generation with Open LVLMs
Analysis Results Table Total Generations
Main 5 prompts (Table 18) Table 1 12,812,400
Main 5 prompts with M1 Table 15 12,812,400
Characteristics prompt with M1,…,M5 Table 7 3,789,000
Characteristics prompt with PATA dataset Table 9 370,050
Additional 6 prompts (Table 19) Table 10 15,374,880
Additional 3 keyword prompts with M1 Table 16 7,687,440
Compute infrastructure.
Licenses of assets used for generation.
F.3 Generation with GPT-4o
Due to cost considerations, we generated responses from GPT-4o using only a subset of the images in SocialCounterfactuals. Specifically, for our study of intersectional gender & physical attributes, we sampled 100 counterfactual sets (containing 10 images each) across 8 occupations (computer programmer, construction worker, doctor, chef, florist, mechanic, chess player, and veterinarian). For race-gender intersectional attributes, we sampled 100 counterfactual sets (containing 12 images each) across 8 occupations (pharmacist, bartender, computer programmer, construction worker, doctor, cashier, dancer, and police officer). For intersectional race & physical attributes, we sampled 35 counterfactual sets (containing 30 images each) across 8 occupations (construction worker, blacksmith, electrician, telemarketer, web developer, software developer, barber, computer programmer). We generate 3 responses for these sampled images to the Characteristics prompt by varying the random seed, producing a total of 78k generations from gpt-4o-2024-05-13 with a maximum token length of 512. We used the API default settings for all other parameters.
Appendix G Lexical Analysis
This section provide additional details and results for the lexical analysis described in Section 4.2.
G.1 Details of the PMI Analysis
[…]
where PMI is calculated as follows:
[…]
G.2 GPT-4 Stereotype Identification
G.3 Additional PMI Stereotype Results
G.4 Stereotype Analysis for GPT-4o
Due to financial constraints, we were unable to perform the entire analysis for GPT-4o. However, we did prompt GPT-4o for 8000 images from the physical-gender subset, 8400 images from the physical-race subset, and 9600 images from the race-gender subset, running the “Characteristics” prompt 3 times for each image. We then performed the PMI analysis (with the parameter min-freq adjusted to 3 to account for the smaller data) and GPT-4 filtering as described above. In general, the GPT-4o results appear to be less problematic than for the other models, suggesting that the guardrails put in place by OpenAI are effective at avoiding stereotypical output. However, this interpretation must be considered in the light of two potentially confounding factors: (1) The data subset is much smaller, and so problems that only surface rarely may not appear in this smaller sample; (2) The filtering step itself involves a GPT-4 model, which may only be sensitive to certain stereotypical words (which it is already programmed to avoid in the generation stage). The results for groups based on physical characteristics and gender are given in Table 24. We observe fewer offensive words relating to obesity and age, and in fact old males are described as authoritative, wise, and computer-savvy. Older females are described in neutral terms which suggest neither positive nor negative views on aging. Table 25 shows the results output for groups defined by physical characteristics and race, and is notably sparse. The only results after the PMI and filtering steps are so-called “positive” stereotypes relating to Asians. Finally, Table 26 shows the results for race-gender intersectional groups. The words are fairly innocuous and mostly related to occupations and descriptive appearance, with the possible exception of associating the adjective modest to Middle Eastern women. Overall, GPT-4o seems less likely to rely on stereotypes when prompted for characteristics of various groups.
Appendix H Stereotype Content Model Analysis for Competency
H.1 Analysis Details
Nicolas et al. [18] present a set of automatically-generated lexicons, based on seed words sourced from the social psychology literature, for a number of different dimensions of stereotype content. This includes warmth (sub-divided into two facets, sociability and morality) and competence (sub-divided into two facets, ability and assertiveness). Words in each lexicon are assigned either a positive (+1) or negative (-1) value according to their direction along that dimension (e.g., the word friendly is associated with positive warmth, while unfriendly is associated with negative warmth, or coldness). We consider the two poles of each dimension separately, leading to four features for each generated text: the number of words associated with competence, the number of words associated with incompetence, the number of words associated with warmth, and the number of words associated with coldness. In the current work, since the dataset is defined by occupation, we focus on competence as being a highly relevant and desirable trait. The normalized counts are computed by dividing the counts for each category by the total number of words in the generated text (after stop-word removal).  We conduct this analysis at the occupation level due to differences in how LVLMs articulate competency-related words across images depciting different occupations. For more accurate estimation, we limit our lexical analyses to occupations for which at least 35 observations were available.
H.2 Additional Results from Analysis of Competency-Associated Words
Figure 13 provides additional results from our lexical analysis of competency-associated words for race-gender and physical-race attribute types. The plots provide the relative proportional representation of intersectional groups which produced the fewest competency-related words, measured across occupations for model responses to the Characteristics prompt. We see similar trends w.r.t. physical characteristics as previously described in Section 4.3; specifically, images depicting obese individuals had the lowest average frequency of competency words across evaluated occupations. For images depicting race-gender attributes, we see lower competency word frequency for Asian, Middle Eastern, Indian, and Latino subjects.
(a) Race-Gender
(b) Physical-Race
Appendix I Numeric Question Analysis
In this experiment, we prompted the LLMs to output numeric responses to prompts, in order to facilitate quantitative comparisons across groups. Because the images all depict occupations, the prompts are focused on job performance. In each case, the prompt specifies the depicted occupation (though not the race, gender, or physical characteristics of the individuals). The five numeric prompts are listed in Table 29. The prompts were generally successful at numeric output that could be interpreted in the context of the question (i.e., a number between 1 and 10, or a number that could be interpreted as a salary); one exception was thee “Salary” prompt for the instructBLIP model, which had to be excluded from the analysis.
For the first four questions, the output should be a number between 1 and 10. If the number was outside that range, it was excluded. In some cases the models embedded the numeric result in a sentence (e.g. The person in the image should be rated as an 8 out of 10.). In such cases, simple NLP techniques were used to extract the relevant number. For the last prompt, the number should be interpreted as an annual salary. If the output text indicated that it was an hourly wage (e.g. I would offer \$12/hr) then the number was multiplied by 2000 to estimate the annual salary. If the salary was less than 1000 or greater than 10,000,000 it was excluded.
@@ Appendix I Numeric Question Analysis
Table 21: PMI results for intersectional groups defined by race and gender.
@@
obese female curvy, caregiver, mechanically-inclined, plus-sized
obese male overweight, plus-sized, heavyset
skinny female []
@@
old female elderly, senior, grey-haired, white-haired, gray-haired
old male elderly, doctor, senior, hardworking, gardener, gardening, computer-savvy, authoritative, wise, handy, scholarly
young female []
young male manual, cook, labor, construction-worker, gamer
tattooed female nurse, working-class
tattooed male player, hipster, urban
Table 24: Analysis of stereotypes in GPT-4o generations on physical-gender subset (“Characteristics” prompt only).
@@
obese Black []
obese Indian []
@@
old Asian hardworking, technologically
old Black []
old Indian []
@@
old Middle Eastern []
@@
skinny Asian helpful, efficient, service-oriented
skinny Black []
skinny Indian []
@@
tattooed Asian []
tattooed Black []
tattooed Indian []
tattooed Latino []

### v2

@@ 8 Ethical Considerations
We believe the findings from this paper will raise awareness of potential risks and harms when LVLMs are deployed at scale. We hope that our work encourages future research aimed at reducing such risks, and believe it will consequently have positive societal impacts through the development of more fair and responsible AI models. Without awareness of the limitations, biases, stereotypes, and toxicity of LVLMs, we risk not just correctness but also fairness to demographic groups. Through our exploration of mitigation strategies in this study, we intend to inspire further innovations for improving fairness in LVLMs.
Appendix A Methodology
In this section, we describe the SocialConterfactuals dataset used in this study, list the open-ended prompts for bias probing in LVLMs, and provide details on output generation with open-source and commercial LVLMs. Further, we provide license information for the used resources and describe the computational infrastructure used for the generation experiments.
A.1 SocialCounterfactuals Dataset
@@
The SocialCounterfactuals dataset was introduced by Howard et al. (2024) and generated automatically using Stable Diffusion with cross-attention control in order to produce visually consistent counterfactual sets. Each set portrays a person in an occupation-based scenario (accountant, pastry chef, plumber, surgeon, etc); however, each image in the set varies in terms of the social attributes of the depicted person. There are three intersectional subsets: Race-Gender, Physical-Gender, and Physical-Race. The size of each subset is given in Table 5. The attribute Gender can have one of two values, male or female. The attribute Race can have one of six values: White, Black, Indian, Asian, Middle Eastern, or Latino. The Physical Characteristics attribute can have one of five values: skinny, obese, young, old, or tattooed.
Subset Counterfactual Images Total
Sets Per Set Images
@@
Table 5: Number of images in each of the three intersectional subsets of SocialCounterfactuals (adapted from Howard et al. (2024)).
A.2 Generating Outputs with LVLMs
Text Prompts:
Table 6 provides the open-ended text prompts investigated in this study, which were adopted from the work by Fraser and Kiritchenko (2024). We refer to each prompt throughout this paper by its short-hand label indicated in Table 6. Table 7 provides additional prompts that we investigated; we provide results for these additional prompts only in the subsequent sections of this Appendix.
Label Prompt
Describe Describe the image in as much detail as possible.
Table 6: Text prompts used for main analyses.
Label Prompt
Table 7: Additional prompts investigated during this study.
Generation with Open LVLMs:
Analysis Results Table Total Generations
Main 5 prompts (Table 6) Table 1 12,812,400
Main 5 prompts with M1 Table 19 12,812,400
Characteristics prompt with M1,…,M5 Table 4 3,789,000
Characteristics prompt with PATA dataset Table 9 370,050
Additional 6 prompts (Table 7) Table 10 15,374,880
Additional 3 keyword prompts with M1 Table 20 7,687,440
Generation with GPT-4o:
Due to cost considerations, we generated responses from GPT-4o using only a subset of the images in SocialCounterfactuals. Specifically, for our study of intersectional gender & physical attributes, we sampled 100 counterfactual sets (containing 10 images each) across 8 occupations (computer programmer, construction worker, doctor, chef, florist, mechanic, chess player, and veterinarian). For race-gender intersectional attributes, we sampled 100 counterfactual sets (containing 12 images each) across 8 occupations (pharmacist, bartender, computer programmer, construction worker, doctor, cashier, dancer, and police officer). For intersectional race & physical attributes, we sampled 35 counterfactual sets (containing 30 images each) across 8 occupations (construction worker, blacksmith, electrician, telemarketer, web developer, software developer, barber, computer programmer). We generate 3 responses for these sampled images to each of our five main prompts by varying the random seed, producing a total of 78k generations per prompt from gpt-4o-2024-05-13 with a maximum token length of 512. We used the API default settings for all other parameters.
A.3 Licenses of Assets Used for Generation
A.4 Compute Infrastructure
Appendix B Differences in Model Response Length across Social Groups
In our experiments, we generate responses of up to 512 tokens from each LVLM. The vast majority of model responses are much shorter than this limit because an end-of-sequence or end-of-turn token terminates generation early. This allows us to study how intersectional social groups depicted in an image influence the amount of text that LVLMs generate in response to different prompts.
@@ Appendix B Differences in Model Response Length across Social Groups
Appendix C Toxicity Analysis
In this section, we provide results for the toxicity probing experiments using a small-scale dataset of real images, PATA. We also conduct a human evaluation of the performance of the Perspective API toxicity detection tool on a sample of generated outputs. Finally, we describe additional experimental results and examples for the toxicity analysis of the LVLMs’ generations on the SocialCounterfactuals dataset.
C.1 Toxicity results for real images from the PATA dataset
Table 9 provides the toxicity scores for different prompts used with images from the real-image dataset PATA. See Section 4.1 for additional discussion of these results.
@@ C.1 Toxicity results for real images from the PATA dataset
C.2 Human Analysis of Automatic Toxicity Evaluation
Due to the scale of LVLM generations which we evaluate in this study, we must rely upon automated methods because it would be infeasible to perform human annotation on over 57 million text sequences. A potential downside of this approach is that automated evaluation methods may themselves possess biases or inaccuracies which could influence our results. Therefore, three of the authors of this work annotated a subset of generations to verify the correctness of automated toxicity classification with the Perspective API.
We randomly sampled 50 LVLM responses to the Personality prompt, divided equally between Race-Gender and Physical-Gender social attribute types and five ranges of toxicity scores which were produced by the Perspective API for the generation: 0-0.2, 0.2-0.4, 0.4-0.6, 0.6-0.8, and 0.8-1.0. Given only the generated text, each annotator labeled the text as either non-toxic/offensive or potentially toxic/offensive. The Fleiss’ Kappa score for the three human annotations on this subset was 0.8788, indicating almost perfect agreement. On this same subset of generations, we found that setting a threshold of 0.3 on the Perspective API toxicity score for binary classification produces nearly perfect agreement with the human annotators, resulting in a Fleiss’ Kappa score of 0.8794 when Perspective API labels are added as a fourth annotator. This shows that human judgements of potentially toxic or offensive content closely match Perspective API evaluations across a broad range of toxicity scores and multiple social attribute types.
Prior studies on the Perspective API have found potential biases in its predictions related to different social groups Sap et al. (2019); Pozzobon et al. (2023). To study whether such biases influence our results, we repeated the same annotation process described above using different set of 50 generations which were sampled only from the subset in which the perceived race of the depicted person is Black and this race is explicitly mentioned by the LVLM in its generation. We find that the three human annotators have a similar level of agreement on this subset, producing a Fleiss’ Kappa of 0.8604. However, agreement with the Perspective API is slightly lower, with a Fleiss’ Kappa of 0.7221. Upon manual inspection of the disagreements, we found that the Perspective API flagged some generations as toxic when the perceived race of the individual is described along with a weapon, such as the generation “…a black woman wearing a police uniform, standing on the side of a busy road with a gun in my holster.” Such cases were not flagged as potentially toxic or offensive by human annotators. Nevertheless, the Fleiss’ Kappa score on this subset still indicates substantial agreement between human annotators and the Perspective API overall.
C.3 Variation in Toxicity Scores by Occupation Depicted in the Image
@@
C.4 MaxToxicity Evaluation with Other Prompts
@@
C.5 Evaluation of Open-Source LVLMs with Other Perspective API Scores
@@
$\textrm{MaxInsult}_{c}=$ $\max_{(a_{i},a_{j})\in A}\bigr{[}I(x)_{c,a_{i},a_{j}}\bigr{]}$ (2)
$-\min_{(a_{i},a_{j})\in A}\bigr{[}I(x)_{c,a_{i},a_{j}}\bigr{]}$
@@
$-\min_{(a_{i},a_{j})\in A}\bigr{[}IA(x)_{c,a_{i},a_{j}}\bigr{]}$
$\textrm{MaxFlirtation}_{c}=$ $\max_{(a_{i},a_{j})\in A}\bigr{[}F(x)_{c,a_{i},a_{j}}\bigr{]}$ (4)
$-\min_{(a_{i},a_{j})\in A}\bigr{[}F(x)_{c,a_{i},a_{j}}\bigr{]}$
@@
To better understand the high values of flirtation scores for BakLLaVA responses to the characteristics prompt (Table 13), we analyze a subset of these generations for 8 occupations which had the highest standard deviation of flirtation scores across physical-gender groups. Figure 7 provides boxplots of BakLLaVA’s Flirtation scores for these responses, broken down by intersectional physical-gender groups. While we observe higher flirtation scores for female subjects in general, skinny, young, and tattooed females have particularly high Flirtation scores relative to other groups across occupations such as Dentist, Bartender, Cashier, and Driver. The high degree of variability across different occupations suggests that bias in the generation of flirtatious content is highly influenced by the occupation depicted in the image.
C.6 Additional Examples
C.7 GPT-4o Refusal Rates
Table 14 provides the refusal percentages of GPT-4o for different physical characteristics and genders. We observed the highest refusal percentage for images depicting obese individuals.
Obese Tattooed Old Young Skinny
Male 4.0% 0.7% 0.6% 0.7% 0.8%
Female 5.9% 1.4% 0.7% 1.1% 1.5%
Appendix D Lexical Analysis of Stereotypes
This section provides additional details and results for the lexical analysis of stereotypes in LVLMs’ generations, described in Section 4.2.
D.1 Details of the PMI Analysis
[…]
where PMI is calculated as follows:
[…]
D.2 Stereotype Identification with GPT-4
To validate this approach to stereotype identification, we conducted a small manual annotation study. We selected the top 50 words (as ranked by PMI) across 6 different intersectional groups, for a total of 300 words. Each word was annotated by 3 annotators as being either stereotypical/offensive for that social group, or not. We then took a majority vote over the three annotations to determine a gold label for each word. When we compare the GPT-4 annotations for the same words and groups, we find a high precision (P = 0.82) but a low recall (R = 0.29). This suggests that the humans are labelling more words as potentially stereotypical/offensive than GPT-4, suggesting that the bias problems reported here are likely a conservative estimate of the situation, and that large-scale human evaluation may uncover even more issues.
D.3 Additional Stereotype Results
D.4 Stereotype Analysis for GPT-4o
Due to financial constraints, we were unable to perform the entire analysis for GPT-4o. However, we did prompt GPT-4o for 8000 images from the physical-gender subset, 8400 images from the physical-race subset, and 9600 images from the race-gender subset, running each prompt 3 times for each image. We then performed the PMI analysis (with the parameter min-freq adjusted to 3 to account for the smaller data) and GPT-4o filtering as described above. In general, the GPT-4o results appear to be somewhat less problematic than other models in terms of the number of identified stereotypes, suggesting that the guardrails put in place by OpenAI are effective at avoiding stereotypical output. However, this interpretation must be considered in the light of two potentially confounding factors: (1) The data subset is much smaller, and so problems that only surface rarely may not appear in this smaller sample; (2) The filtering step itself involves a GPT-4o model, which may only be sensitive to certain stereotypical words (which it is already programmed to avoid in the generation stage). The results for groups based on physical characteristics and gender are given in Table 24. We observe fewer outright offensive words relating to obesity and age, although obese males are still described as hunched, misunderstood, shy, and associated with bullying, and older men and women are described as slow and slowing. Table 25 shows the results output for groups defined by physical characteristics and race. Twelve groups have no results after the PMI and filtering steps are applied. The images of tattooed subjects yield some of the most stereotypical results, with both tattooed Black and Latino subjects associated with words like gang(s), underworld, criminal, illegal, and warfare. While white tattooed individuals are associated with some similar terms, we also see evidence for an alternative interpretation of tattoos as a vehicle of self-expression in words like: individualistic, edgy, rogue, and alternative. Finally, Table 26 shows the results for race-gender intersectional groups. In addition to the stereotypes of Black males mentioned in main part of the paper, we do observe some problematic associations (Middle Eastern men with bandits, Middle Eastern women with harem, and Indian men with poverty and corruption). However, many of the terms also appear to be more cultural than stereotypical in nature.
Appendix E Lexical Analysis of Competency
This section provides additional details and results for the lexical analysis of competency-associated words in LVLMs’ generations, described in Section 4.3.
E.1 Analysis Details
Nicolas et al. (2021) present a set of automatically-generated lexicons, based on seed words sourced from the social psychology literature, for a number of different dimensions of stereotype content. This includes warmth (sub-divided into two facets, sociability and morality) and competence (sub-divided into two facets, ability and assertiveness). Words in each lexicon are assigned either a positive (+1) or negative (-1) value according to their direction along that dimension (e.g., the word friendly is associated with positive warmth, while unfriendly is associated with negative warmth, or coldness). We consider the two poles of each dimension separately, leading to four features for each generated text: the number of words associated with competence, the number of words associated with incompetence, the number of words associated with warmth, and the number of words associated with coldness. In the current work, since the dataset is defined by occupation, we focus on competence as being a highly relevant and desirable trait. The normalized counts are computed by dividing the counts for each category by the total number of words in the generated text (after stop-word removal).  We conduct this analysis at the occupation level due to differences in how LVLMs articulate competency-related words across images depicting different occupations. For more accurate estimation, we limit our lexical analyses to occupations for which at least 35 observations were available.
E.2 Additional Results from Analysis of Competency-Associated Words
Figure 12 shows the relative proportional representation of intersectional groups that produced the fewest competency-related words, measured across occupations for model responses to the Characteristics prompt, normalized by length of each generated sequence. Across all models, images depicting obese and old individuals produced the fewest number of competency-associated words. For images depicting race-gender attributes, we see lower competency word frequency for Asian, Middle Eastern, Indian, and Latino subjects.
(a) Physical-Gender
(b) Race-Gender
(c) Physical-Race
Appendix F Numeric Question Analysis
Label Prompt
Table 15: Prompts used to generate numerical responses.
This section provides additional details and results for the numeric question analysis described in Section 4.4. In this experiment, we prompted the LLMs to output numeric responses to prompts, in order to facilitate quantitative comparisons across groups. Because the images all depict occupations, the prompts are focused on job performance. In each case, the prompt specifies the depicted occupation (though not the race, gender, or physical characteristics of the individuals). The five numeric prompts are listed in Table 15. The prompts were generally successful at numeric output that could be interpreted in the context of the question (i.e., a number between 1 and 10, or a number that could be interpreted as a salary); one exception was the “Salary” prompt for the instructBLIP model, which had to be excluded from the analysis.
For the first four questions, the output should be a number between 1 and 10. If the number was outside that range, it was excluded. In some cases the models embedded the numeric result in a sentence (e.g., The person in the image should be rated as an 8 out of 10.). In such cases, simple NLP techniques were used to extract the relevant number. For the last prompt, the number should be interpreted as an annual salary. If the output text indicated that it was an hourly wage (e.g., I would offer \$12/hr) then the number was multiplied by 2000 to estimate the annual salary. If the salary was less than 1000 or greater than 10,000,000 it was excluded.
Appendix G LLM-LVLM Mapping
In Section 5.1, we analyze the relationship between MaxToxicity observed in LVLMs and that of the LLM from which they were trained. Table 16 provides a mapping between the LVLMs used in our study and the LLMs from which they were derived.
LVLM LLM
LLaVA-7B Vicuna-7b-v1.5
LLaVA-13B Vicuna-13b-v1.5
BakLLaVA Mistral-7B-Instruct-v0.1
InstructBLIP Vicuna-7b-v1.5
LLaVA-Gemma Gemma-2b-it
Table 16: Mapping of LVLMs to the LLM from which they were trained
Appendix H Inference-Time Mitigation
@@
M1 M2 M3 M4 M5
Model Mean 90% Mean 90% Mean 90% Mean 90% Mean 90%
[…]
@@
Table 17 provides details of the five mitigation instructions that we investigated. Mitigation instructions are added either before or after our original prompts, which are represented with the {Prompt} placeholder. Table 18 provides full results for the MaxToxicity reduction achieved by the M1,…, M5 mitigation instructions for physical-gender images and the Characteristics prompt, including both the mean and 90th percentile reductions. In addition to evaluating mitigation effectiveness for physical-gender images and the Characteristics prompt (Table 4), we also evaluated the effectiveness of the M1 mitigation instruction using all three intersectional attribute types and other prompts used throughout our analyses. Tables 19 and 20 provide the reduction in MaxToxicity scores when the M1 mitigation instruction is used across these various evaluation settings. We find that this instruction is most effective for the Personality and Trustworthiness prompts, particularly for images depicting intersectional physical-gender social attributes. However, the high degree of variability across prompts (even for the same model and social attribute types) indicates that a single inference-time mitigation instruction is unlikely to reduce bias across a broad-range of generation scenarios. In many cases (e.g., BakLLaVA physical-gender and physical-race evaluations), large reductions in MaxToxicity for the Personality, Characteristics, and Trustworthiness prompts are contrasted with increases in MaxToxicity for the Describe, Backstory, and Pretend prompts. This highlights the need for additional research into robust methodologies for reducing bias in LVLMs.
@@
obese female curvy, loose
obese male stocky, overweight, heavier, boomer, shorts, football, bullying, burger, diet, giant, loneliness, joke, introverted, hunched, misunderstood, exercise, gaming, size, sturdy, shy, anxiety, chicken, fried, eating
skinny female blonde, female
@@
old female geriatric, geriatrics, age-related, widow, elderly, seniors, elder, aging, retiring, seniority, white-haired, grey-haired, mature, retirement, aged, graying, slow, old-fashioned, retire, senior
old male widower, thinning, mature, grey-haired, graying, twilight, white-haired, seniority, rusty, gray-haired, retiring, balding, fifties, middle-aged, retirement, maturity, elderly, venerable, mid-fifties, sixties, slowing, bald, mid-forties, slower, veteran, senior, elder, mid-thirties, aged, older, old-fashioned
young female fragile, cleaning, dolls, blouse, baking
young male []
tattooed female rebellious, corrupt, vigilante, prejudice, criminal, stereotypes
tattooed male gang, rebellious, edgy, tribal, nonconformist, martial, piercings, violence, rebellion, at-risk, rough, graffiti, troubled, gritty, toughness, punk, gangs, crime, violent, stereotypical, stigma, biker, poverty, illegal, tough, underworld, criminal, prejudices, prejudice, marginalized, homeless, struggles, struggling, blue-collar, stereotypes
Table 24: Analysis of stereotypes in GPT-4o generations on physical-gender subset.
@@
obese Black afro, basketball, saxophone, hip-hop, jazz
obese Indian low-cost, middle-class, rural, high-paying, second-hand
@@
old Asian wrinkles, wrinkled, elderly, bonsai, mature, tea, graying, calligraphy, ceremonies, temples, maturity, kendo, gardening, samurai, experienced, medicine, sage, fishing, fisherman, antique, serene, martial, contemplative, aging, old-fashioned
old Black afro, jazz
old Indian turban, homemaker, sitar, agricultural, elderly, farmer, retiring, wrapped, mustache, pottery, agriculture, rural, village, headscarf, laborer, farm, labor-intensive, manual, villagers, seamstress, labor, weaver
@@
old Middle Eastern headscarf, turban, desert, dust
@@
skinny Asian shanghai, china, kyoto, japan, tokyo, osaka, neo-tokyo, kendo, culinary, immigrants, martial
skinny Black basketball, afro, hip-hop, jazz
skinny Indian baskets, sitar, cricket, yoga, rural, agricultural, homemaker, middle-class, villages, meditation, farmer, underdeveloped, low-cost, immigrant
@@
tattooed Asian irezumi, samurai, katana, martial, kendo, dragon
tattooed Black oppressed, gangs, corrupt, underworld, intimidating, impoverished, fighting, vigilante, graffiti, hacker, rebellious, enforcement, underbelly, corruption, illegal, authorities, criminal, shadowy, clandestine, hacking, fight, crime, shadow, tough, street, warfare
tattooed Indian tribal, tribe
tattooed Latino rebellious, notoriety, street, gang, rebellion, troubled, underbelly, illegal, vigilante, underworld